# RefVNLI: Towards Scalable Evaluation of Subject-driven Text-to-image Generation

### v1

@@ 8 Reproducibility
Misalignment Injection Instructions (Short Captions) 1. Understand the Caption: Carefully read the short caption to fully grasp the scene it describes. 2. Identify and Swap: Select a single visual detail within the caption to modify. Replace this detail with a different, incorrect, but still plausible visual detail. For example, you might change a color, an object, or a location. Do not modify the underlined entity (if any). 3. Apply the Tags: Enclose the original visual detail within <swap> tags. Immediately after the closing </swap> tag, write the new, incorrect visual detail. There should be no space between the closing </swap> and the new word. Example: If the original sentence is ”The cat sat on the red mat,” and you want to change ”red” to ”blue,” the result should be: ”The cat sat on the <swap>red</swap><blue> mat.” 4. Final Check: Ensure the modified caption is grammatically correct and reads naturally, even though it now contains a factual error. The sentence should be internally logical, despite contradicting the actual visual content. Again, ensure the underlined entity (if any) remains completely unchanged.
@@
9 Further Information on the Data Construction Pipeline
9.1 Collection of Subject-driven Image Pairs
When collecting the subject-driven image pairs, we added several filtering operations to reduce noise. These included removing blurred images, as well as those not portraying the subject, fur which we used a Gemini [58]. Additionally, for the dataset sourced from TVQA+, we also filtered images containing subtitles and credits, also using Gemini.
9.2 Generation of prompt-imagetgt Hard-negatives
@@
10 Additional Qualitative Examples for Subject Preservation Evaluation
In Fig. 8, Fig. 9, Fig. 10, and Fig. 11 we present additional qualitative examples of subject preservation evaluation for the Animal, Human, Object, and Landmark categories, respectively.

### v2

@@ Appendix B Additional Details on the Meta-evaluation
Misalignment Injection Instructions (Short Captions) 1. Understand the Caption: Carefully read the short caption to fully grasp the scene it describes. 2. Identify and Swap: Select a single visual detail within the caption to modify. Replace this detail with a different, incorrect, but still plausible visual detail. For example, you might change a color, an object, or a location. Do not modify the underlined entity (if any). 3. Apply the Tags: Enclose the original visual detail within <swap> tags. Immediately after the closing </swap> tag, write the new, incorrect visual detail. There should be no space between the closing </swap> and the new word. Example: If the original sentence is "The cat sat on the red mat," and you want to change "red" to "blue," the result should be: "The cat sat on the <swap>red</swap><blue> mat." 4. Final Check: Ensure the modified caption is grammatically correct and reads naturally, even though it now contains a factual error. The sentence should be internally logical, despite contradicting the actual visual content. Again, ensure the underlined entity (if any) remains completely unchanged.
@@
Appendix C Further Information on the Data Construction Pipeline
C.1 Collection of Subject-driven Image Pairs
To reduce noise when collecting subject-driven image pairs, we applied several filtering steps, including the removal of blurred images and those not depicting the intended subject. Subject presence was verified using Gemini (Team, 2024) (version gemini-2.0-flash). For the subset sourced from TVQA+, we additionally filtered out frames containing subtitles or credits, also using Gemini.
For identity-sensitive image pairs, we used Stable Diffusion (Rombach et al., 2022) for inpainting, [^7] https://huggingface.co/stabilityai/stable-diffusion-2-inpainting with $\eta=1.0$ and a guidance scale of 3.0. We retained only images with a full mask size of at least 60,000 pixels (20,000 for humans, focusing on facial regions). Five patches of 250–300 pixels were randomly sampled and inpainted. To increase the likelihood of meaningful subject changes, we further filtered out inpainted images where all patchwise MSE values fell below 6,500 for objects, 5,400 for animals, and 20,000 for humans.
C.2 Collection of Image-prompt Pairs
For the image-captioning of the imagetgt with the inserted bounding boxes, we employed Gemini (Team, 2024) (version gemini-2.0-flash).
C.3 Generation of prompt-imagetgt Hard-negatives
@@
Textual Alignment Subject Preservation Unified Evaluation
RefVNLI (ours) 81.5 87.7 / 86.3 97.0 82.7 83.0 / 62.8 82.3 82.1 85.3 / 72.7 89.0
      reverse classification order 80.0 85.2 / 85.5 95.3 80.9 84.3 / 68.7 87.0 80.4 84.7 / 76.2 91.0
      multiclass 79.5 83.7 / 84.7 94.7 79.6 76.0 / 61.1 86.3 79.5 79.7 / 71.0 90.3
      no markup 78.4 87.0 / 84.3 92.3 65.5 75.9 / 60.8 88.7 71.4 81.1 / 70.6 90.5
      only video-based training 80.4 85.2 / 83.8 96.3 77.6 77.5 / 66.4 89.7 79.0 81.2 / 74.1 92.9
Table 5: Ablation Study: ROC AUC scores for various ablated versions of RefVNLI across benchmarks (over all subjects). The ablations evaluate alternative output formulations, such as reversed classification order, a four-label multiclass framework, and separate aspect classification via a special token. Additional variants exclude subject markup from input prompt, merge reference and target images, or remove identity-sensitive training examples by using only video-based instances.
Appendix D Ablations
To assess the impact of various design decisions in RefVNLI, we run an ablation study examining alternative input and output configurations. On the output side, we test: reversing the classification order (subject preservation before textual alignment); a 4-label multiclass framework for joint text-image alignment classification; and a model that prefixes a designated token (‘TEXT’ or ‘IMAGE’) to the prompt to enable separate classification of each aspect within a unified model. For inputs, we explore the effect of removing subject markup from the prompt and of concatenating imageref and imagetgt instead of passing them separately. Finally, we also explore the impact of omitting the identity-sensitive training examples (see §2.1), by only using the video-based instances during training.
Results (Table 5) show that reversing the classification order degrades performance, particularly in subject preservation, as does evaluating each aspect separately. This suggests that first evaluating textual alignment helps in subject preservation assessment. The multiclass approach also underperforms compared to our dual binary classification setup, highlighting the benefits of treating each criterion independently. Further, removing subject markup weakens subject preservation, underscoring its role in linking the reference image to the prompt. Additionally, concatenating images instead of processing them separately harms performance, emphasizing the advantage of distinct image inputs. Finally, excluding identity-sensitive training instances leads to notable drops in subject preservation, underscoring their importance.
DreamBench++ ImagenHub KITTEN
Animal Human Object Animal Object Multi-subj. Landmarks
OWL-ViT 77.4 87.1 81.8 74.8 83.1 52.0 85.6
GroundingDINO 76.9 85.6 83.4 77.6 84.1 56.8 90.2
Table 6: ROC AUC scores of Crop-IR for subject preservation when employed with two different object-detection models: OWL-ViT and GroundingDino. Bold indicates the highest score per column.
Appendix E Crop-IR Object Detection Model Ablation
Deploying the Crop-IR metric (Winter et al., 2024) requires an object detection model to locate and crop the referenced subjects. To this end, we compare two prominent object detection models: OWL-ViT (Minderer et al., 2022) [^8] https://huggingface.co/google/owlv2-base-patch16 and GroundingDino (Liu et al., 2024a). [^9] https://huggingface.co/IDEA-Research/grounding-dino-base As shown in Table 6, GroundingDino leads to better evaluation of subject preservation, and is therefore adopted to ensure a fairer evaluation.
Inference Time (seconds) GPU Memory Usage (GiB) API Calls Cost (\$)
CLIP 0.1 1.2 -
DINO 0.06 0.7 -
Crop-IR 0.6 5.8 -
ArcFace 1.2 - -
CLIPScore 0.07 0.6 -
BLIPScore 0.7 4.5 -
SigLIP 0.03 1.4 -
TIFA 22.5 26.4 -
VQAScore 0.2 23.1 -
VIEScore 6.9 - 0.04
DreamBench++ (text) 3.2 - 0.02
DreamBench++ (ref) 1.0 - 0.03
[…]
PaliGemmaref 0.4 12.5 -
RefVNLI 0.5 12.5 -
Table 7: Computational costs for all baseline models and RefVNLI, including per-instance inference time (in seconds), GPU memory usage (in GiB), and GPT-4o API costs (in \$, only when applicable), averaged across benchmarks. For DreamBench++, we report separate values for each evaluation criterion, as each requires a distinct API call under its framework. The final three rows present models fine-tuned on our dataset, with PaliGemmatext and PaliGemmaref being the variants tuned exclusively for evaluating textual alignment and subject preservation, respectively.
Appendix F Computational Cost
Table 7 presents the computational costs of all baseline models and RefVNLI, in terms of inference time, GPU memory usage and GPT-4o API costs.
Appendix G Additional Qualitative Examples for Subject Preservation Evaluation
In Figures 9, 10, 11, and 12 we present additional qualitative examples of subject preservation evaluation for the Animal, Human, Object, and Landmark categories, respectively.
@@ Appendix G Additional Qualitative Examples for Subject Preservation Evaluation
Appendix H List of Data and Software Licenses Employed in this Paper
Our framework dependencies are:
  1. 1.
    Mementos dataset: https://github.com/si0wang/Mementos, Misc.
  2. 2.
    TVQA+ dataset: https://github.com/jayleicn/TVQAplus/blob/master/LICENSE, under the MIT License.
  3. 3.
    Open Images dataset: https://github.com/openimages/dataset/blob/main/LICENSE, under an Apache License 2.0.
  4. 4.
    PaliGemma model: https://ai.google.dev/gemma/terms, under Gemma Terms of Use License.
  5. 5.
    Gemini model: https://ai.google.dev/gemini-api/docs/models, under an Apache License 2.0.
  6. 6.
    GPT-4o model: https://github.com/openai/openai-openapi/blob/master/LICENSE, under the MIT License.
  7. 7.
    OWL-ViT model: https://huggingface.co/google/owlv2-base-patch16, under an Apache License 2.0.
  8. 8.
    GroundingDino model: https://huggingface.co/IDEA-Research/grounding-dino-base, under an Apache License 2.0.
  9. 9.
    Stable Diffusion inpainting model: https://huggingface.co/stabilityai/stable-diffusion-xl-base-1.0/blob/main/LICENSE.md, under Stability AI CreativeML Open RAIL++-M License.